OpenStudy (anonymous):

use elmination method to solve system of equations a+8b=17 a+3b=2

OpenStudy (anonymous):

because the variable a already have the same coefficients you can just subtract the two equations (a+8b)-(a+3b)=17-2 a+8b-a+3b=15 11b=15 b=15/11

OpenStudy (anonymous):

then just substitute b with 15/11 to solve for a

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

how did you get 17

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

17 is the answer of the first equation a+8b=17 2 is the answer of the second equation a+3b=2

OpenStudy (anonymous):

so the subtration is used to elminate

OpenStudy (anonymous):

yes, it the equations were a+8b=17 -a+3b=2 then you would use addition to eliminate the a

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

how do I get 17 in the equation when replacing 15/11 for b

OpenStudy (anonymous):

a+8(15/11)=17 a=17-120/11 a=67/11

OpenStudy (anonymous):

am confused because for my answer choices it has a. one solution type ordered pair b. many solutions c. no solutions

OpenStudy (anonymous):

wait that's wrong, i think i did my math wrong

OpenStudy (anonymous):

ahh i see, i subtracted wrong, when you subtract it should be (a+8b)-(a+3b) a+8b-a-3b 5b=17-2=15 b=3

OpenStudy (anonymous):

substitute b with 3 in one of the original equations a+3(3)=2 a+9=2 a=-7 check to see if it works in the first equation -7+8(3)=17 -7+24=17 17=17

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

yea :)